OpenStudy (courtney_celiene1236):

Write the sum using summation notation, assuming the suggested pattern continues. -9 - 3 + 3 + 9 + ... + 81

OpenStudy (sshayer):

tn=a+(n-1)d tn=81 a=-9 d=9-3=6 find d then sum \[=\frac{ n }{ 2 }\left( a+tn \right)\]

OpenStudy (mathmale):

Is this an arithmetic sequence or series, or a geometric sequence or series? What is the common difference?

OpenStudy (courtney_celiene1236):

Isn't it a geometric sequence? And the common difference is 6.

OpenStudy (sshayer):

it is an arithmetic sequence and common difference d=6

OpenStudy (courtney_celiene1236):

Oh. So would i set that up with the Sigma sign?

OpenStudy (courtney_celiene1236):

Would it be this? \[\sum_{n=0}^{15}(-9+6n)\]

OpenStudy (courtney_celiene1236):

@sleepyjess could you help me with this?